Assume the price of a good increase from $6 to $8, leading to a fall in quantity demanded from 50 to 40 units. Calculate the price elasticity of demand for the good at this price range and explain how total revenue will be impacted by the increase in price?
1
Expert's answer
2016-05-27T10:19:03-0400
by definition, the price elasticity of demand is:
e = (dQ/Q) / (dP/P), therefore:
e = (-10/50) / (2/6) = -0.6;
-1 < e = -0.6 < 0; the demand for the good is inelastic

Total Revenue = Price x Quantity Demanded
before the increase in price:
Total Revenue1 = $6 x 50 = $300
after:
Total Revenue2 = $8 x 40 = $320

As we may see from the above figures, the rise in price for the good led to the increased revenue. That happened because of the inelastic demand when the 33% rise in price ( dP/P = $2/$6 = 0.33 = 33%) caused demand to contract by only 20% (dQ/Q = 10/50 = 0.2 = 20%).